Problem 27-6 (Algo) Suppose that in 1984 the total output in a single-good economy was 10,000 buckets of chicken. Also assumeIn 2005, real GDP =
Answer #1

ANSWER:

A) GDP price index = (price of chicken in 1984 / price of chicken in 2005) * 100 = (15 / 20) * 100 = 0.75 * 100 = 75

B) Price level change = (price of chicken in 2005 / price of chicken in 1984) * 100 - 1 = (20 / 15) - 1 = 1.3333 - 1 = 0.3333 or 33.33%

C) Real gdp in 1984 = price of chicken in 2005 * quantity of chicken in 1984 = $20 * 10,000 = $200,000

Real gdp in 2005 = price of chicken in 2005 * quantity of chicken in 2005 = $20 * 22,000 = $440,000